While driving in his car on a snowy day, Mr. Gaglardi slams on the brakes at the bottom of a hill. He then slides with an initial velocity of 14.50m/s [up the slope]. If the acceleration of the car is 4.25m/s2 [down the slope],


a) How far up the slope does the car slide before coming to a momentary rest?

b) If the car continues to accelerate, at what time will it return to a point 2.50m up the slope from its original starting point? (Note that it will be sliding back down at this point!)


1
Expert's answer
2021-03-02T18:05:10-0500

a)


"v^2=2as\\14.5^2=2(4.25)s\\s=24.74\ m"

b)


"y=2.5=14.5t-(0.5)4.25t^2\\t=6.65\ s"